OpenStudy (anonymous):

log base 7 + log base x =2

OpenStudy (anonymous):

Log(base 7) (X)=2 --------------------------------------… Log(base 7) (X)=2 the rule is: ... Base raised to (the answer which is the exponent) = the # we're taking the log of Therefore: (base 7) raised to the exponent of 2 = the # "x" 7 ^4= x (7)(7)(7)(7) = x (49)(49) = x 2401 = x ---------------------------------- CHECK: x = 2,401 Log(base 7) (X) = 4 Log(base 7) (2,401) = 4 ...use the Change of Base Formula to check the answer and see if it's right Log(base 10) of the # divided by Log(base 10) of the base Log(base 10) of 2,401 divided by Log(base 10) of 7 3.38039216 / 0.84509804 = 4 4 = 4 ...TRUE Therefore x = 2,401 ------------------------------------ FINAL ANSWER: x = 2, 401

OpenStudy (anonymous):

rohangrr-- the correct answer is 100/7. I just wanted to know how to get there....thanks!!! You see, I have the answers to the equations, but I have to know how to work them.